OpenStudy (anonymous):

A rock falls from a tower that is 400 ft high. As it is falling its height is given by the function h(t)=400-16t^2. Ho many seconds will it take for the rock to hit the ground (h=0)?

OpenStudy (anonymous):

set \[h(t)=400-16t^2=0\] solve for t

OpenStudy (anonymous):

since h =0 when it touches the ground 400-16t^2=0 16t^2=400 t^2=400/16 t^2= 25 t=5 seconds

OpenStudy (anonymous):

\[0=400-16t^2\] \[16t^2=400\] \[t^2=25\] \[t=5\]

OpenStudy (anonymous):

oh. i see thank you